You are here: Foswiki>ARGOS Web>Commissioning>DiaryRun11>DayTime20160311 (07 Apr 2016, TommasoMazzoni)Edit Attach

SX QC and Pyramid co-pointing

We switch on the AGW and we configure it to be used with IS. FW1: 600-1000, we center on QC.

Board -72.3 -50.1 35.0
Then we offset the TipTilt mirror by [2.2; 2.1]V, corresponding to the position measured in August 2015 during the QC alignment.

We measure the range of FW2 positions that do not vignette the QC beam: -0.02 - 0.13 filter positions. Instead the range of positions for FW2 that are able to feed the QC are: -0.015 - 0.015.

Since FW2 has 6 filters (60deg/filter) we get that 1milli filter corresponds to about 1mrad of tilt. The distance between QC and FW2 is about 150mm, so 1mfilter corresponds to about 80mas on-sky. Considering a positioning accuracy of 3mfilter for the FW2 this corresponds to 150mas of co-pointing error on-sky due to FW2 histeresys. A lambda/D on the Pyramid tip corresponds to about 30mas. The co-pointing error due to FW2 misplacement is about 5 lambda/D, comparable with the modulation diameter...

We scan various LUCI rotator positions and we check the modulation offset that allow to co-point PYR and QC. Moreover we check if moving FW2 we can minimize the co-pointing offset:

Rot PI FW2 Notes
345 [1.6; 4.4] -0.002  
[1.7; 4.7] [0.0]  
[0.5; >5.0] 0.003  
[2.2; 2.1] -0.007  
255 [2.2; 2.1] 0.003 we've lost 1/3 of light wrt -0.007
165 [0.6; 3.0] -0.007 we cannot co-point with [2.2; 2.1]
75 [-0.80; 0.3] -0.007 we cannot co-point with [2.2; 2.1]
We notice that flux on QC is increasing with FW2 position going from + to -: for example we have 100% flux at -0.007, 125% at -0.015 and 50% at 0.005. This effect is independent of rotator position.

We test the QC in closed loop from AARB w/o disturb. Co-point in truth sensing seems to work. Working at 75deg of LUCI rotator angle we have a co-pointing offset of [-0.5, 0.5]V.

We test the QC also with 1" equivalent disturb, using old (December) rec:

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20160311_191232 0.5,0.4,0.3 93.5948 20151213_063700 -1 T F
Since the ASM is skipping when 1" disturb is applied we try to reduce the alignment coma using the haxapod. After optimization:

SX Hexa 4.65 1.77 -0.71 8 -107
Board -75 -48 45    
We close again with 1" disturb:

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20160311_193515 2,0.45,0.45 105.398 20160310_213600 -1 T F
20160311_193631 2,0.45,0.45 102.52 20160310_213600 -1 T F
20160311_193651 2,0.45,0.45 106.676 20160310_213600 -1 T F
Since co-pointing can work between -15 and 165deg of LUCI rotator we try to close the loop on QC in these positions, checking the co-pointing convergence:

Rot Board PI
75 [-75; 48] [2.2; 2.1]
30 [-74.5; -50] [1.3; 1.6]
0 [-74.4; - 51.5] [2.3; 2.0]
-15 [-74.5; -52.5] [3.0; 2.1]
165 [-79.4; -49.6] [0.77; -3.0]
WBoardPositionAndRotatorAngle.png

While closing loop at different rotator angles we also acquired snapshots:

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20160311_193753 2,0.45,0.45 98.1175 20160310_213600 -1 T F 75deg
20160311_193923 2,0.45,0.45 98.7126 20160310_213600 -1 T F
20160311_213447 1,0.4,0.3 68.7564 20160310_213600 -1 T F 0deg
20160311_213516 1,0.4,0.3 64.1665 20160310_213600 -1 T F
20160311_214225 0.2,0.1,0.05 78.831 20160310_213600 -1 T F -15deg
20160311_214248 0.2,0.1,0.05 84.0563 20160310_213600 -1 T F
20160311_215056 1,0.4,0.3 58.2939 20160310_213600 -1 T F 165deg
20160311_215114 1,0.4,0.3 59.6251 20160310_213600 -1 T F
Then we close with a 1" disturb, source is equivalent to mag10 star:

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20160311_215502 1,0.4,0.3 72.7737 20160310_213600 -1 T F
20160311_215815 1,0.4,0.3 75.9512 20160310_213600 -1 T F
20160311_215910 1,0.4,0.3 75.5551 20160310_213600 -1 T F
20160311_220025 1,0.4,0.3 77.5041 20160310_213600 -1 T F
Then we reduce the flux to test the truth sensig at magnitude 16:

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20160311_225330 0.5,0.3,0.1 63.924 20160310_213600 -1 T F w/o dist
20160311_225359 0.5,0.3,0.1 65.7635 20160310_213600 -1 T F
20160311_225531 1,0.4,0.4 83.072 20160310_213600 -1 T F 1" dist
20160311_225551 1,0.4,0.4 85.7034 20160310_213600 -1 T F
20160311_225619 1,0.4,0.4 86.4308 20160310_213600 -1 T F

WhiteLightSource SX to Pyramid magnitude calibration

We scan the white light source aperture and we check the equivalent magnitude on the pyramid:

WLS steps 303 80 70 65 60 55 50 45 40
Pyr mag (600-1000) - 10 11 11.5 13.5 15 16 17 18
Pyr mag (90-10) 11.5 12 13 13.5 15        

QC hot spot on CCD47

We checked the CCD47 pixel corresponding to QC tip: FW1: silver mirror (#4), FW2: empty (#1) -> [34; 42] @ BIN16.

LgswUnitDx to CalUnit alignment

We aligned the LgswUnitDx to the calibration unit. We ended up to a limit position were the decenter Y of the hexapod was hitting the limit; we have less coma with respect to yesterday, but not null

DX Hexa -5.95 -7.6 0.95 145 -20 0
We feed all the three Lgsw and we acquired a 10 modes interaction matrix with KL_v21

IM REC COMBO
20160311_204825 20160311_205800 20160311_205900
We closed the AO loop without disturbance with the jitter initial position 20160311_210436 and the pupil motors in :

BEAM X Y
BLUE 415 515
YELLOW 205 370
RED 435 345

LgswUnitDx FoV measurement

We do a sweep of +/-6" in 11 steps and we measured the FOV center of the BLUE

20160311_214300_fov_center.png

We do a sweep of +/-6" in 11 steps around the RED FOV center estimated from the previous image and we measured the FOV center of the RED

20160311_220000_fov_center.png

We updated the procedure in order to obtain a squared grid of the hexapod movements. We did a sweep of +/-8" in 7 steps around the YELLOW FOV center estimated from the first image and we measured the FOV center of the YELLOW;

20160311_230300_fov_center.png

BEAM [Rx; Ry]
Blue [138.6; -20.1]
Yellow [142.12; -19.16]
Red [147; -24]
We defined a new subapertures set to align the center of the FOV to the center of the subapertures and minimize the jitter offset. The new set is 20160311_234700.

Since we cannot shift the BLUE subapertures in the x direction, we change the pointing moving the hexapod The new BLUE hexapod position is

DX Hexa -5.95 -7.6 0.95 141 -22 1.41
We also changed the BLUE pupil motor position

BEAM X Y
BLUE 410 530
YELLOW 205 370
RED 435 345

PockelsCellsTransmission test

Side PCT
DX 20160312_005745_pcTransmission.jpg
SX 20160312_055924_pcTransmission.jpg

TipTiltSystemSx IM recording

We recorded a new IM for the SX QC using the AGW IS.

Board -72.8 -49.9 25

  • IM: 20160312_030023

The scan seems not very good, it is not passing by the center of the QC, so we refine the source centering on the QC and acquire a new IM. QC speed is 500Hz, Pyramid (TS) speed is 100Hz.

Board -72.82 -49.85 25

IM REC COMBO REC
20160312_030807 20160312_030900 20160312_032600
We change FW1 to 600-1000 to be able to ramp up the IS and acquire the IM at 1kHz: 20160312_032258. But this new one is worse because we have a double spot on QC.

-- TommasoMazzoni - 12 Mar 2016
Topic revision: r4 - 07 Apr 2016, TommasoMazzoni
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback